# Dipak Meher    Carlotta Domeniconi

### v1

@@ Abstract
Human smuggling networks are increasingly adaptive and difficult to analyze. Legal case documents offer critical insights but are often unstructured, lexically dense, and filled with ambiguous or shifting references, which pose significant challenges for automated knowledge graph (KG) construction. While recent LLM-based approaches improve over static templates, they still generate noisy, fragmented graphs with duplicate nodes due to the absence of guided extraction and coreference resolution. The recently proposed CORE-KG framework addresses these limitations by integrating a type-aware coreference module and domain-guided structured prompts, significantly reducing node duplication and legal noise. In this work, we present a systematic ablation study of CORE-KG to quantify the individual contributions of its two key components. Our results show that removing coreference resolution results in a 28.32% increase in node duplication and a 4.32% increase in noisy nodes, while removing structured prompts leads to a 4.34% increase in node duplication and a 73.33% increase in noisy nodes. These findings offer empirical insights for designing robust LLM-based pipelines for extracting structured representations from complex legal texts.
@@ I Introduction
Human smuggling has become a complex operation involving dynamic networks of actors, routes, vehicles, and intermediaries [1]. These networks facilitate illicit human mobility for payment, exploiting legal loopholes and adapting rapidly to enforcement efforts to avoid detection. Often linked to transnational crime, they pose serious security and humanitarian risks due to their covert, profit-driven nature. Understanding their structure is critical for shaping effective policy responses and disrupting illicit operations. However, much of the relevant intelligence remains locked within unstructured legal texts such as court rulings and case transcripts, making systematic analysis challenging.
@@
Prior work has shown the value of knowledge graphs in legal investigations. Mazepa et al. [2] and Shi et al. [3], for instance, used rule-based methods to build graphs from legal cases, but their reliance on static templates limits flexibility in handling aliasing and surface-level entity variation. Large language models (LLMs) have shown promise in knowledge extraction [4], yet their use in human smuggling case narratives remains limited. A major challenge is merging semantically equivalent mentions of typed entities in complex legal texts [5], which, if left unresolved, creates fragmented and redundant graphs. Additionally, LLM-based extractors often misclassify or hallucinate entities and relations [6], adding noise that hinders downstream analysis.
@@
To address these challenges, we introduced CORE-KG [7], a modular framework that builds clean and interpretable knowledge graphs from legal case documents using large language models. CORE-KG combines two main components: (1) a type-aware coreference resolution module that sequentially resolves contextually similar mentions within each entity type, and (2) a structured prompting strategy that guides the model to extract relevant entities and relationships while filtering legal boilerplate and reducing ambiguity. Compared to a GraphRAG-based baseline, CORE-KG reduces node duplication by 33.28% and legal noise by 38.37%, showing strong improvements in graph quality.
@@
In this paper, we conduct an ablation study on CORE-KG to quantify the individual contributions of its two key components: type-aware coreference resolution and structured prompting. In the first setting (CoreKG-no-coref), we disable the coreference module while keeping structured prompts to assess its role in entity consolidation. In the second (CoreKG-no-structprompts), we retain coreference resolution but remove structured prompts to evaluate their impact on reducing noise and improving extraction precision. We adopt the same setup as CORE-KG, using legal case documents related to human smuggling from publicly available U.S. federal and state court proceedings. Since no gold-standard graphs exist, we compare the outputs of each ablation with CORE-KG and a GraphRAG-based baseline [8], both of which generate knowledge graphs from raw legal text using large language models.
@@
Our ablation study reveals that removing either coreference resolution or structured prompts significantly degrades KG quality. Specifically, removing coreference resolution (CoreKG-no-coref) results in a 28.32% increase in node duplication and a 4.32% increase in noisy nodes. In contrast, removing structured prompts (CoreKG-no-structprompts) leads to a 73.33% increase in noisy nodes and a 4.34% increase in node duplication. On average, this corresponds to a 38.83% degradation in noise reduction and a 16.33% degradation in entity resolution, underscoring the necessity of both modules for accurate and interpretable knowledge graph construction. [^1] Our code is available at https://github.com/dipakmeher/corekg-ablation-study.git.
@@
  • We quantitatively show that removing the coreference module results in a 28.32% increase in node duplication and 4.32% increase in noise, while removing structured prompting leads to a 4.34% increase in duplication and a 73.33% spike in noise, demonstrating the distinct and complementary impact of each component.
@@ II Related Work
Knowledge graphs enable structured representation of unstructured text for tasks like reasoning, retrieval, and analysis [8]. They have seen wide adoption across domains including education [9], life sciences [10], and construction safety [11]. However, poorly constructed graphs often exhibit node duplication and structural fragmentation, which hinder their effectiveness [12].
@@ II-A Knowledge Graph Construction
Traditionally, knowledge graphs are constructed by extracting key entities and their relationships from text and linking them sequentially. Entity extraction relies on rule-based, statistical, and domain-specific methods [13], while relation extraction uses syntactic, lexical, and semantic features [14], ontology-based systems like HowNet [15], and semi-supervised techniques [16]. Though effective in narrow domains, these approaches often rely on hand-crafted rules and focus on maintaining close alignment with the source text.
More recently, LLM-based frameworks now enable prompt-driven knowledge graph construction with minimal manual effort [6, 17]. Kommineni et al. [18] combined competency question generation, ontology design, and RAG-based triple extraction, while Zhang et al. [6] proposed a modular pipeline with open IE, schema definition, and canonicalization. However, these methods often assume clean input and overlook challenges like reference ambiguity and entity aliasing, which are common in legal texts where entities shift across aliases, roles, and pronouns. To address these, coreference resolution is critical for maintaining graph coherence. Wang et al. [19] demonstrated that even sparse pronoun usage can impact graph completeness, underscoring the need to integrate coreference mechanisms into the construction pipeline.
Several works have explored knowledge graph construction in criminal domains. Mazepa et al. [2] used rule-based NLP pipelines and CoreNLP to build a homicide investigation graph. Shi et al. [3] employed regex-based extraction to construct a Neo4j graph for job-related crime indictments. While effective in legal contexts, both rely on static templates and lack coreference handling and modular prompting, limiting their robustness in complex narratives.
@@ II-B Coreference Resolution
Several methods have been proposed to address node duplication and fragmentation via coreference resolution [20]. CNN-based models by Pogorilyy et al. [21] and Wu et al. [22] enhance semantic and syntactic pattern modeling.
Recent advances in large language models (LLMs) have advanced coreference resolution in low-resource settings, using prompt-based, few/zero-shot, and chain-of-thought strategies [23, 24]. However, these methods often target single entity types or general narratives without modeling type distinctions [23, 25]. While effective in focused contexts, their ability to generalize across diverse, role-shifting entity types in domain-specific texts remains underexplored.
Some progress has been made in coreference resolution for legal texts. Jia et al. [5] introduced a neural model combining ELMo, BiLSTM, and GCNs to resolve speaker-based coreference in court records. However, to the best of our knowledge, no prior work has examined coreference resolution in the legal domain using large language models. While instruction-tuned LLMs have shown competitive performance in coreference resolution [4], these studies are restricted to standard narrative data and do not address the challenges of multi-entity legal documents. This raises an open question: to what extent can LLM-based coreference resolution scale to complex, multi-entity legal texts, and how much does it contribute to improving knowledge graph quality?
@@ III-A1 Coreference Resolution
We design a Coreference Resolution module that leverages the contextual reasoning capabilities of large language models (LLMs) [26, 27]. The model unifies disparate mentions under a canonical form (e.g., “Young” and “the defendant”), ensuring that all relational links converge on a single entity node. This improves structural consistency and enhances analytical clarity.
However, resolving all entity types simultaneously can dilute the model’s attention and cause type drift (incorrect category assignments) or feature entanglement (interference between features of different types) [28, 29]. For instance, the phrase “The Camp” might denote a temporary migrant holding site (Location) but may be misclassified as an Organization in ambiguous contexts. Similarly, “the van” might refer to either a transport vehicle (Means of Transportation) or a holding site (Location).
@@
    The legal text is first passed to the LLM with a Person-focused prompt, linking mentions like “Young,” “the defendant,” and “the driver” to a canonical form.
@@ III-A2 Entity-Relationship Extraction and KG Construction
The coreference-resolved text is fed into the Knowledge Graph Construction (KGC) module of the GraphRAG framework [8], which comprises: (1) a KGC component that extracts entities and relationships to form a structured graph, and (2) a retrieval module that uses this graph as an index for response generation. This work focuses on the KGC component, extended for processing legal case files on human smuggling.
@@
(1) Sequential Entity Extraction to Reduce Attention Spread: GraphRAG’s default joint extraction can dilute model attention across entity types, leading to missed entities, misclassifications, or loss of fine-grained distinctions, issues common in legal texts with overlapping entity mentions [30]. To address this, we enforce a fixed extraction order: Person first, followed by Location, Routes, Organization, etc. The model extracts relationships only after completing all entity types in sequence. This reduces attention competition, improves entity precision, and ensures more complete and reliable knowledge graph construction.
@@
(3) Entity Type Definitions to Reduce Misclassification: LLMs often misclassify entities due to overgeneralization e.g., labeling event names as Location just because they appear near geographic terms during training [31, 32]. This problem worsens in legal texts, where subtle context matters. To reduce such errors, we include clear definitions and examples for all seven entity types directly in the prompt. These definitions help the model distinguish between similar types like Person, Organization, and Route, leading to more accurate and context-aware extractions.
@@ III-B Ablation Study: Isolating the Impact of Coreference Resolution and Structured Prompts
We construct three experimental variants for comparison. The first is the Full CORE-KG pipeline, which includes both coreference resolution and structured prompts. The second variant, denoted as CoreKG-no-coref, removes the coreference module while retaining the structured prompt for entity and relationship extraction. In this setting, raw legal text is passed directly to the KGC module without resolving intra-document entity references. The third variant, CoreKG-no-structprompt, removes our structured extraction prompt and instead uses the default prompt from GraphRAG [8], while still applying coreference resolution. This allows us to isolate the contribution of structured prompt design while holding the input (coref-resolved text) constant.
@@ IV-C Baselines and Experimental Design
To evaluate the individual contributions of the two core modules in the CORE-KG pipeline—coreference resolution and structured prompting—we design an ablation study comparing four system variants:
  • GraphRAG: A minimally adapted version of the GraphRAG framework [8], using its standard prompt. It is applied directly to the input text, without coreference resolution or modified structured prompts.
@@ IV-D Evaluation Measures
In addition to these metrics, we perform a quality-to-analysis comparison by examining the actual graph structures across CORE-KG, CoreKG-no-coref, and CoreKG-no-str-prompt. This graph-level analysis provides qualitative insight into the nature of duplication, noise, connectivity, and semantic coherence, offering a deeper understanding of how different components impact overall graph quality.
@@ V-A1 Node Duplication
In CoreKG-no-coref, the use of structured prompts alone reduces duplication compared to GraphRAG, as shown in Table I. Specifically, the average node duplication rate decreases from 30.53% in GraphRAG to 26.01% in CoreKG-no-coref. This improvement stems from in-prompt filtering, sequential extraction steps, and the explicit definition of the seven entity types described in the Method section relevant to human smuggling analysis.
However, structured prompts cannot resolve referential variations on their own. Coreference resolution remains essential for merging such mentions. For instance, in Case 6 (Figure 2(b)), “Interstate Highway 35,” “Interstate,” and “I-35” all refer to the same route. Without coreference resolution, these are treated as distinct entities, as shown in Figure 2(b). This results in a higher node duplication rate in CoreKG-no-coref (22.92%) compared to CORE-KG (11.90%).
@@ V-A2 Noisy Nodes
The removal of coreference resolution leads to a moderate increase in noisy nodes, rising from 16.65% in CORE-KG to 17.36% in CoreKG-no-coref, corresponding to a relative degradation of 4.26% (Table II). This shows that coreference resolution helps reduce noise, though its effect is smaller than on duplication. It refines the input text by consolidating scattered or semantically equivalent mentions that might otherwise increase text density, diffuse the LLM’s attention, and result in noisier extractions or misclassifications. While the improvement is modest, it contributes to overall graph cleanliness.
Even without coreference resolution, CoreKG-no-coref achieves comparatively low noise levels. This indicates that the structured prompt itself plays a key role in reducing noise by guiding the LLM to extract only relevant entity and relationship types. Through in-prompt filtering, sequential extraction, and predefined entity types tailored to human smuggling analysis, the structured prompt suppresses irrelevant or overly generic mentions. The sharp rise in noise from CoreKG-no-coref to GraphRAG (57.84%) further underscores the value of these structured instructions.

### v2

@@ Abstract
Human smuggling networks are increasingly adaptive and difficult to analyze. Legal case documents offer critical insights but are often unstructured, lexically dense, and filled with ambiguous or shifting references, which pose significant challenges for automated knowledge graph (KG) construction. While recent LLM-based approaches improve over static templates, they still generate noisy, fragmented graphs with duplicate nodes due to the absence of guided extraction and coreference resolution. The recently proposed CORE-KG framework addresses these limitations by integrating a type-aware coreference module and domain-guided structured prompts, significantly reducing node duplication and legal noise. In this work, we present a systematic ablation study of CORE-KG to quantify the individual contributions of its two key components. Our results show that removing coreference resolution results in a 28.25% increase in node duplication and a 4.32% increase in noisy nodes, while removing structured prompts leads to a 4.29% increase in node duplication and a 73.33% increase in noisy nodes. These findings offer empirical insights for designing robust LLM-based pipelines for extracting structured representations from complex legal texts.
@@ I Introduction
Human smuggling has become a complex operation involving dynamic networks of actors, routes, vehicles, and intermediaries [carrasco2025scapegoating]. These networks facilitate illicit human mobility for payment, exploiting legal loopholes and adapting rapidly to enforcement efforts to avoid detection. Often linked to transnational crime, they pose serious security and humanitarian risks due to their covert, profit-driven nature. Understanding their structure is critical for shaping effective policy responses and disrupting illicit operations. However, much of the relevant intelligence remains locked within unstructured legal texts such as court rulings and case transcripts, making systematic analysis challenging.
@@
Prior work has shown the value of knowledge graphs in legal investigations. Mazepa et al. [mazepa2022relationships] and Shi et al. [shi2022knowledge], for instance, used rule-based methods to build graphs from legal cases, but their reliance on static templates limits flexibility in handling aliasing and surface-level entity variation. Large language models (LLMs) have shown promise in knowledge extraction [le2023large], yet their use in human smuggling case narratives remains limited. A major challenge is merging semantically equivalent mentions of typed entities in complex legal texts [ji2020deep], which, if left unresolved, creates fragmented and redundant graphs. Additionally, LLM-based extractors often misclassify or hallucinate entities and relations [zhang2024extract], adding noise that hinders downstream analysis.
@@
To address these challenges, we introduced CORE-KG [meher2025llm], a modular framework that builds clean and interpretable knowledge graphs from legal case documents using large language models. CORE-KG combines two main components: (1) a type-aware coreference resolution module that sequentially resolves contextually similar mentions within each entity type, and (2) a structured prompting strategy that guides the model to extract relevant entities and relationships while filtering legal boilerplate and reducing ambiguity. Compared to a GraphRAG-based baseline, CORE-KG reduces node duplication by 33.28% and legal noise by 38.37%, showing strong improvements in graph quality.
@@
In this paper, we conduct an ablation study on CORE-KG to quantify the individual contributions of its two key components: type-aware coreference resolution and structured prompting. In the first setting (CoreKG-no-coref), we disable the coreference module while keeping structured prompts to assess its role in entity consolidation. In the second (CoreKG-no-structprompts), we retain coreference resolution but remove structured prompts to evaluate their impact on reducing noise and improving extraction precision. We adopt the same setup as CORE-KG, using legal case documents related to human smuggling from publicly available U.S. federal and state court proceedings. Since no gold-standard graphs exist, we compare the outputs of each ablation with CORE-KG and a GraphRAG-based baseline [edge2024local], both of which generate knowledge graphs from raw legal text using large language models.
@@
Our ablation study reveals that removing either coreference resolution or structured prompts significantly degrades KG quality. Specifically, removing coreference resolution (CoreKG-no-coref) results in a 28.25% increase in node duplication and a 4.32% increase in noisy nodes. In contrast, removing structured prompts (CoreKG-no-structprompts) leads to a 73.33% increase in noisy nodes and a 4.29% increase in node duplication. On average, this corresponds to a 38.83% increase in noise and a 16.27% increase in node duplication, underscoring the necessity of both modules for producing accurate and interpretable knowledge graphs. [^1] Our code is available at https://github.com/dipakmeher/corekg-ablation-study.git.
@@
  • We quantitatively show that removing the coreference module results in a 28.25% increase in node duplication and 4.32% increase in noise, while removing structured prompting leads to a 4.29% increase in duplication and a 73.33% spike in noise, demonstrating the distinct and complementary impact of each component.
@@ II Related Work
Knowledge graphs enable structured representation of unstructured text for tasks like reasoning, retrieval, and analysis [edge2024local]. They have seen wide adoption across domains including education [chen2018knowedu], life sciences [callahan2024open], and construction safety [fang2020knowledge]. However, poorly constructed graphs often exhibit node duplication and structural fragmentation, which hinder their effectiveness [huaman2020duplication].
@@ II-A Knowledge Graph Construction
Traditionally, knowledge graphs are constructed by extracting key entities and their relationships from text and linking them sequentially. Entity extraction relies on rule-based, statistical, and domain-specific methods [sun2018overview], while relation extraction uses syntactic, lexical, and semantic features [kambhatla2004combining], ontology-based systems like HowNet [liu2007implementation], and semi-supervised techniques [carlson2010coupled]. Though effective in narrow domains, these approaches often rely on hand-crafted rules and focus on maintaining close alignment with the source text.
More recently, LLM-based frameworks now enable prompt-driven knowledge graph construction with minimal manual effort [zhang2024extract, vizcarra2024representing]. Kommineni et al. [kommineni2024human] combined competency question generation, ontology design, and RAG-based triple extraction, while Zhang et al. [zhang2024extract] proposed a modular pipeline with open IE, schema definition, and canonicalization. However, these methods often assume clean input and overlook challenges like reference ambiguity and entity aliasing, which are common in legal texts where entities shift across aliases, roles, and pronouns. To address these, coreference resolution is critical for maintaining graph coherence. Wang et al. [wang2020coreference] demonstrated that even sparse pronoun usage can impact graph completeness, underscoring the need to integrate coreference mechanisms into the construction pipeline.
Several works have explored knowledge graph construction in criminal domains. Mazepa et al. [mazepa2022relationships] used rule-based NLP pipelines and CoreNLP to build a homicide investigation graph. Shi et al. [shi2022knowledge] employed regex-based extraction to construct a Neo4j graph for job-related crime indictments. While effective in legal contexts, both rely on static templates and lack coreference handling and modular prompting, limiting their robustness in complex narratives.
@@ II-B Coreference Resolution
Several methods have been proposed to address node duplication and fragmentation via coreference resolution [liu2023brief]. CNN-based models by Pogorilyy et al. [pogorilyy2019coreference] and Wu et al. [wu2017deep] enhance semantic and syntactic pattern modeling.
Recent advances in large language models (LLMs) have advanced coreference resolution in low-resource settings, using prompt-based, few/zero-shot, and chain-of-thought strategies [das2024co, gan2024assessing]. However, these methods often target single entity types or general narratives without modeling type distinctions [das2024co, tran2025coreference]. While effective in focused contexts, their ability to generalize across diverse, role-shifting entity types in domain-specific texts remains underexplored.
Some progress has been made in coreference resolution for legal texts. Jia et al. [ji2020deep] introduced a neural model combining ELMo, BiLSTM, and GCNs to resolve speaker-based coreference in court records. However, to the best of our knowledge, no prior work has examined coreference resolution in the legal domain using large language models. While instruction-tuned LLMs have shown competitive performance in coreference resolution [le2023large], these studies are restricted to standard narrative data and do not address the challenges of multi-entity legal documents. This raises an open question: to what extent can LLM-based coreference resolution scale to complex, multi-entity legal texts, and how much does it contribute to improving knowledge graph quality?
@@ III-A1 Coreference Resolution
We design a Coreference Resolution module that leverages the contextual reasoning capabilities of large language models (LLMs) [wang2023can, havrilla2024glore]. The model unifies disparate mentions under a canonical form (e.g., “Young” and “the defendant”), ensuring that all relational links converge on a single entity node. This improves structural consistency and enhances analytical clarity.
However, resolving all entity types simultaneously can dilute the model’s attention and cause type drift (incorrect category assignments) or feature entanglement (interference between features of different types) [zhou2023universalner, abdelnabi2024you]. For instance, the phrase “The Camp” might denote a temporary migrant holding site (Location) but may be misclassified as an Organization in ambiguous contexts. Similarly, “the van” might refer to either a transport vehicle (Means of Transportation) or a holding site (Location).
@@
    The legal text is first passed to the LLM with a prompt focused type Person, linking mentions like “Young,” “the defendant,” and “the driver” to a canonical form.
@@ III-A2 Entity-Relationship Extraction and KG Construction
The coreference-resolved text is fed into the Knowledge Graph Construction (KGC) module of the GraphRAG framework [edge2024local], which comprises: (1) a KGC component that extracts entities and relationships to form a structured graph, and (2) a retrieval module that uses this graph as an index for response generation. This work focuses on the KGC component, extended for processing legal case files on human smuggling.
@@
(1) Sequential Entity Extraction to Reduce Attention Spread: GraphRAG’s default joint extraction can dilute model attention across entity types, leading to missed entities, misclassifications, or loss of fine-grained distinctions, issues common in legal texts with overlapping entity mentions [abdelnabi2025get]. To address this, we enforce a fixed extraction order: Person first, followed by Location, Routes, Organization, etc. The model extracts relationships only after completing all entity types in sequence. This reduces attention competition, improves entity precision, and ensures more complete and reliable knowledge graph construction.
@@
(3) Entity Type Definitions to Reduce Misclassification: LLMs often misclassify entities due to overgeneralization e.g., labeling event names as Location just because they appear near geographic terms during training [peters2025generalization, dai2024bias]. This problem worsens in legal texts, where subtle context matters. To reduce such errors, we include clear definitions and examples for all seven entity types directly in the prompt. These definitions help the model distinguish between similar types like Person, Organization, and Route, leading to more accurate and context-aware extractions.
@@ III-B Ablation Study: Isolating the Impact of Coreference Resolution and Structured Prompts
We construct three experimental variants for comparison. The first is the Full CORE-KG pipeline, which includes both coreference resolution and structured prompts. The second variant, denoted as CoreKG-no-coref, removes the coreference module while retaining the structured prompt for entity and relationship extraction. In this setting, raw legal text is passed directly to the KGC module without resolving intra-document entity references. The third variant, CoreKG-no-structprompt, removes our structured extraction prompt and instead uses the default prompt from GraphRAG [edge2024local], while still applying coreference resolution. This allows us to isolate the contribution of structured prompt design while holding the input (coref-resolved text) constant.
@@ IV-C Baselines and Experimental Design
To evaluate the individual contributions of the two core modules in the CORE-KG pipeline, coreference resolution and structured prompting, we conduct an ablation study comparing four system variants:
  • GraphRAG: A minimally adapted version of the GraphRAG framework [edge2024local], using its standard prompt. It is applied directly to the input text, without coreference resolution or modified structured prompts.
@@ IV-D Evaluation Measures
In addition to these metrics, we perform a graph-level comparison across CORE-KG, CoreKG-no-coref, and CoreKG-no-str-prompt. This analysis provides qualitative insight into the nature of duplication, noise, relational cohesion, and semantic coherence, offering a deeper understanding of how different components impact overall graph quality.
@@ V-A1 Node Duplication
In CoreKG-no-coref, using structured prompts alone already reduces duplication compared to GraphRAG, lowering the average node-duplication rate from 30.53% to 26.01%, as shown in Table II. This reduction comes primarily from the in-prompt filtering rules, sequential extraction steps, and the explicit definition of entity types relevant to human smuggling analysis.
However, structured prompts alone cannot resolve referential variations across the text. As shown in the CoreKG-no-coref graph for Case 6 in Figure 2(b), the mentions “Interstate Highway 35,” “Interstate,” and “I-35” all refer to the same route. Without coreference resolution, these surface variations are extracted as separate nodes, fragmenting what should be a single entity. This leads to a higher node-duplication rate in CoreKG-no-coref (26.01%) compared to the full CORE-KG pipeline (20.28%).
@@ V-A2 Noisy Nodes
The removal of coreference resolution leads to a moderate increase in noisy nodes, rising from 16.65% in CORE-KG to 17.37% in CoreKG-no-coref, a relative degradation of 4.32% (Table II). This shows that coreference resolution helps reduce noise, though its effect is smaller than on duplication. It refines the input text by consolidating scattered or semantically equivalent mentions that might otherwise increase text density, diffuse the LLM’s attention, and result in noisier extractions or misclassifications. While the improvement is modest, it supports a clearer and more accurate graph.
Even without coreference resolution, CoreKG-no-coref maintains comparatively low noise levels. This indicates that the structured prompt plays a key role in reducing noise by guiding the LLM to extract relevant and precise entities and relationships. Through in-prompt filtering, sequential extraction, and clearly defined entity-type categories, the structured prompt suppresses irrelevant or overly generic mentions. The sharp rise in noise from CoreKG-no-coref (17.37%) to GraphRAG (27.43%) highlights the importance of structured prompting, since GraphRAG does not include these additional prompt-level instructions, resulting in poorer performance.